# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 2.3. Cancellation Lemma
$\mathcal{J}_{l,p}^{1}[g](v)$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{-l}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\frac{(1-\lambda_{v-\nu})^{1-l/2}}{\lambda_{v-\nu}}\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}\frac{|f(v-\nu)|^{p}}{|v-\nu|^{2}}\mathds{1}_{|v-\nu|>|v|/2>2}\,d\nu$
$\approx{\langle}v{\rangle}^{-2}\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\,d\nu={\langle}v{\rangle}^{-2}\|f\|_{L_{v}^{p}}^{p}.$
Estimate (2.26) is proved.
Estimate for $\mathcal{J}_{l,p}^{2}$
Fix $v\in{\mathbb{R}}^{3}$. We use (2.20) and (2.7) to bound
$\mathcal{J}_{l,p}^{2}[f](v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v_{1}^{*})|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v_{1}^{*}|>|v|/2>2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{-}(u))|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{\left\langle v-R_{\sigma}^{-}(u)\right\rangle^{l}\left(1-\lambda_{v-R_{\sigma}^{+}(u)}|\widehat{(v-R_{\sigma}^{-}(u))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(u)|>|v|/2>2}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{-}(y))|^{p}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|^{2}}{{\langle}v-R_{\sigma}^{-}(y){\rangle}^{l}\left(1-\lambda_{v-R_{\sigma}^{-}(y)}|\widehat{(v-R_{\sigma}^{-}(y))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(y)|>|v|/2>2}\,dy\,d\sigma.$
Using Proposition 2.1 to substitute $\nu:=R_{\sigma}^{-}(y)$, we obtain
$\mathcal{J}_{l,p}^{2}[f](v)\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{\alpha}{\langle}v-\nu{\rangle}^{-l\alpha}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l\alpha/2}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{\alpha}{\langle}v-\nu{\rangle}^{-l\alpha}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\int_{{\mathbb{S}}^{2}}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l\alpha/2}}\,d\sigma\,d\nu.$
Following an identical argument to the one used for $\mathcal{J}_{l,p}^{1}$, we conclude (2.26). ∎
@@ 2.4. Pre-post collisional change of variables
(2.27) $\begin{cases}v^{*}=\frac{v+v_{1}}{2}-\frac{|v-v_{1}|}{2}\sigma\\ v_{1}^{*}=\frac{v+v_{1}}{2}+\frac{|v-v_{1}|}{2}\sigma\\ \eta=\frac{v_{1}-v}{|v-v_{1}|}\end{cases}$
@@
Second, writing $u^{*}:=v^{*}-v_{1}^{*}$, we compute
$u^{*}\cdot\eta=-|v-v_{1}|\frac{\sigma\cdot(v_{1}-v)}{|v-v_{1}|}=u\cdot\sigma,$
@@ 3. Convolution estimates for Maxwell molecules
In the section, we prove the convolution estimates for Maxwell molecules. Given $b$ satisfying (H.1)–(H.3), we define the Maxwell collisional operator with angular cross-section $b$ as
(3.1) $Q_{M,b}^{+}(f,g)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Slightly abbreviating notation we will be writing $Q^{+}_{M}:=Q^{+}_{M,1}$ for the Maxwell operator with constant cross-section equal to $1$.
As mentioned, when $b$ is constant, the previous works [1, 2] provide an estimate only when $r<3$. Here, we are able to extend the estimates to arbitrary $r$, as long as $p<3$.
3.1. The $L_{v}^{\infty}$– estimate for Maxwell molecules
First, we prove the convolution estimate corresponding to $r=\infty$.
@@ 3.1. The $L_{v}^{\infty}$– estimate for Maxwell molecules
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{\infty}_{v}$ satisfying the estimate
@@
For $p=1$, $q=\infty$, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{1}_{v}\times C_{b}^{1}({\mathbb{R}}^{3})\to L^{\infty}_{v}$ satisfying the estimate
(3.3) $\|Q^{+}_{M,b}\|_{L^{\infty}_{v}}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}},$
where $C_{b}^{1}({\mathbb{R}}^{3})$ denotes the space of continuously differentiable and bounded functions in $v$.
@@
We start by proving the first claim. By density, it suffices to show (3.2) for $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Since $p<3$, we have $\frac{1}{2q}<\frac{1}{p}$. Pick $\alpha>1/2$ with $\alpha p<q$. We use (H.3) and Hölder’s inequality to bound
@@
Now for the case $p=1$, $q=\infty$, again by density, it suffices to show (3.3) for $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Assume $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$ and $g\in C_{b}^{1}({\mathbb{R}}^{3})$. By (H.3) and (2.18), we bound
(3.5) $|Q^{+}_{M,b}(f,g)|$ $\leq\|g\|_{L^{\infty}_{v}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}},$
@@
3.2. The $L_{v}^{r}$– estimate for Maxwell molecules
Now, we prove the convolution estimate for general $1\leq r<\infty$.
@@ 3.2. The $L_{v}^{r}$– estimate for Maxwell molecules
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q,r<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{r}_{v}$ satisfying the estimate
@@
By density, it suffices to prove (3.6) for $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. We take different cases for $1<p<3$ and $p=1$.
@@
Take $\alpha>1/2$ with $\frac{\alpha}{p^{\prime}}<\frac{1}{q^{\prime}}+\frac{1}{r}$, which is possible due to the above bound. Equivalently, we can write this relation as $1-\frac{p^{\prime}}{\alpha q^{\prime}}<\frac{p^{\prime}}{\alpha r}$. Pick $s\in{\mathbb{R}}$ with $1-\frac{p^{\prime}}{\alpha q^{\prime}}<s<\frac{p^{\prime}}{\alpha r}$. Then $\frac{(1-s)\alpha q^{\prime}}{p^{\prime}}<1$ and $\frac{s\alpha r}{p^{\prime}}<1$.
@@
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha(1-s)q^{\prime}/p^{\prime}}}\,d\sigma\,dv_{1}\,dv,$
@@
Raising the above bound to the $r$ and integrating in $v$, and using (2.28), we obtain
@@
(3.10) $\leq I_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\right)^{1/r},$
@@
Raising the above expression to the $r$-th power and integrating in $v$, we obtain
@@ 4. Convolution estimates for hard-spheres
In this final section, we prove our main result, Theorem 1.1. As mentioned in Subsection 1.2, we decompose $Q^{+}$ into two pieces $Q_{0},\,Q_{1}$ corresponding to the outgoing particle with the largest velocity, and estimate each piece separately. This is the content of Propositions 4.14.4 which are the heart of our contributions. Upon establishing those estimates, the claim of Theorem 1.1 follows easily.
Namely, given $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$, conservation of energy implies that either $|v^{*}|^{2}\geq E/2$ or $|v_{1}^{*}|^{2}\geq E/2$. Thus, using (H.3), we can bound
(4.1) $|Q^{+}(f,g)|\lesssim|Q_{0}(f,g)|+|Q_{1}(f,g)|,$
@@
(4.2) $Q_{0}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1},$
@@
(4.3) $Q_{1}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
In the following, we prove convolution estimates for the operators $Q_{0},Q_{1}$.
4.1. The $L_{v}^{\infty}$– estimate for $Q_{0}$
Proposition 4.1.
Let $0<\gamma\leq 1$. Consider $1\leq p_{0},q_{0}\leq\infty$ with $p_{0}\leq 2$, such that $\frac{1}{p_{0}}+\frac{1}{q_{0}}=1$. Then for any $k\geq l>2$, there holds the estimate
(4.4) $\|{\langle}v{\rangle}^{k}Q_{0}(f,g)\|_{L_{v}^{\infty}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p_{0}}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q_{0}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{k}:={\langle}v{\rangle}^{k}f$, $g_{l}:={\langle}v{\rangle}^{l}g$. By the triangle inequality, we can bound $Q_{0}$ as follows
(4.5) $|Q_{0}(f,g)|\lesssim|Q_{0}^{sm}(f,g)|+|Q_{0}^{r}(f,g)|+|Q_{0}^{s}(f,g)|,$

### v4

@@ 2.3. Cancellation Lemma
$\mathcal{J}_{l,p}^{1}[g](v)$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{2-l}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\frac{(1-\lambda_{v-\nu})^{1-l/2}}{\lambda_{v-\nu}}\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}\frac{{\langle}v-\nu{\rangle}^{2}}{|v-\nu|^{2}}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|g(v-\nu)|^{p}\,d\nu=\|g\|_{L_{v}^{p}}^{p}.$
Estimate (2.26) is proved. ∎
@@ 2.4. Pre-post collisional change of variables
(2.27) $\begin{cases}v^{*}=\frac{v+v_{1}}{2}-\frac{|v-v_{1}|}{2}\sigma\\ v_{1}^{*}=\frac{v+v_{1}}{2}+\frac{|v-v_{1}|}{2}\sigma\\ \eta=-\widehat{u}=\frac{v_{1}-v}{|v-v_{1}|}\end{cases}$
@@
Second, writing $u^{*}:=v^{*}-v_{1}^{*}=-|u|\sigma$, we compute
$u^{*}\cdot\eta=-|u|\sigma\cdot(-\widehat{u})=u\cdot\sigma,$
@@ 3. Convolution estimates for Maxwell molecules
In the section, we prove the convolution estimates for Maxwell molecules. Given $b$ satisfying (H.1)–(H.3), we define the Maxwell bilinear collisional operator with angular cross-section $b$ as
(3.1) $Q_{M,b}^{+}(f,g)=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\mathds{1}_{\widehat{u}\cdot\sigma>0}\,d\sigma\,dv_{1}.$
We note that, when $f=g$, (H.2) implies
$Q_{M,b}^{+}(f,f)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\,d\sigma\,dv_{1},$
thus (3.1) extends (1.9) for $\gamma=0$. Slightly abbreviating notation we will be writing $Q^{+}_{M}:=Q^{+}_{M,1}$ for the Maxwell operator with constant cross-section equal to $1$.
As mentioned, when $b$ is constant, the previous works [1, 2] only provide an estimate when $r<3$. Here, we are able to extend the estimates to arbitrary $r$, as long as $p<3$.
3.1. The $L_{v}^{\infty}$– estimate
First, we prove the convolution estimates corresponding to $r=\infty$.
@@ 3.1. The $L_{v}^{\infty}$– estimate
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then, $Q_{M,b}^{+}$ extends uniquely to a bounded, bilinear operator $Q_{M,b}^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{\infty}_{v}$ satisfying the estimate
@@
For $p=1$, $q=\infty$, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{1}_{v}\times C_{b}({\mathbb{R}}^{3})\to L^{\infty}_{v}$ satisfying the estimate
(3.3) $\|Q^{+}_{M,b}\|_{L^{\infty}_{v}}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}}.$
@@
We start by proving the first claim. By density, it suffices to show (3.2) for $f,g\in C_{c}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}({\mathbb{R}}^{3})$. Since $p<3$, we have $\frac{1}{2q}<\frac{1}{p}$. Hence we can choose $\alpha>1/2$ such that $\alpha p<q$. We use (H.3) and Hölder’s inequality to bound
@@
Now for the case $p=1$, $q=\infty$, again by density, it suffices to show (3.3) for $f\in C_{c}({\mathbb{R}}^{3})$. Assume $f\in C_{c}({\mathbb{R}}^{3})$ and $g\in C_{b}({\mathbb{R}}^{3})$. By (H.3) and (2.18), we bound
(3.5) $|Q^{+}_{M,b}(f,g)|$ $\lesssim\|g\|_{L^{\infty}_{v}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}},$
@@
3.2. The $L_{v}^{r}$– estimate
Now, we prove the convolution estimates for $1\leq r<\infty$.
@@ 3.2. The $L_{v}^{r}$– estimate
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q,r<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$. Then, $Q_{M,b}^{+}$ extends uniquely to a bounded, bilinear operator $Q_{M,b}^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{r}_{v}$ satisfying the estimate
@@
By density, it suffices to prove (3.6) for $f,g\in C_{c}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}({\mathbb{R}}^{3})$. We treat the cases $1<p<3$ and $p=1$ separately.
@@
Hence we can choose $\alpha>1/2$ such that $\frac{\alpha}{p^{\prime}}<\frac{1}{q^{\prime}}+\frac{1}{r}$. Equivalently, we can write this condition as $1-\frac{p^{\prime}}{\alpha q^{\prime}}<\frac{p^{\prime}}{\alpha r}$. Pick $s\in{\mathbb{R}}$ with $1-\frac{p^{\prime}}{\alpha q^{\prime}}<s<\frac{p^{\prime}}{\alpha r}$. Then $\frac{(1-s)\alpha q^{\prime}}{p^{\prime}}<1$ and $\frac{s\alpha r}{p^{\prime}}<1$.
@@
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha(1-s)q^{\prime}/p^{\prime}}}\,d\sigma\,dv_{1}\,dv,$
@@
(3.10) $\leq I_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r},$
@@
Raising the expression above to the $r$-th power and integrating in $v$, we obtain
@@ 4. Convolution estimates for hard-spheres
In this final section, we prove our main result, Theorem 1.1, for the bilinear operator $Q^{+}(f,g)$ given in (1.13). We first establish the necessary estimates for smooth, compactly supported functions for $r=\infty$ and $2<r<\infty$ respectively.
4.1. The $L_{v}^{\infty}$– estimate
Proposition 4.1.
Let $0<\gamma\leq 1$ and $b$ satisfying (H.1)–(H.3). Consider $1<p<q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$ and $k\geq l>1$ with $k\geq\gamma+\frac{2}{p}$, $l>\frac{2}{p}$. Then, there holds the estimate
(4.1) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in C_{c}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>1$ with $k\geq\gamma+\frac{2}{p}$ and $l>2/p$. Fix $f,g\in C_{c}({\mathbb{R}}^{3})$ and denote $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$.
We use the triangle inequality to bound $Q^{+}$:
(4.2) $|Q^{+}(f,g)|\lesssim|Q^{sm}(f,g)|+|Q^{r}(f,g)|+|Q^{s_{1}}(f,g)|+|Q_{1}^{s_{2}}(f,g)|,$
@@ 4.1. The $L_{v}^{\infty}$– estimate
$Q_{sm}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|\leq 4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1},$
$Q_{r}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{|v^{*}|\geq|v|/2}\,d\sigma\,dv_{1},$
$Q_{s,1}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
$Q_{s,2}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1},$
and given $v\in{\mathbb{R}}^{3}$, we denoted
(4.3) $\mathcal{B}_{1}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v|}{2},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|\geq{\langle}v{\rangle}^{-1/2}\right\},$
(4.4) $\mathcal{B}_{2}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v|}{2},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<{\langle}v{\rangle}^{-1/2}\right\}.$
Estimate for $Q_{sm}^{+}$
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k+1-\gamma}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v|\leq 4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim E^{\gamma/2}(1+E)^{-k/2}\lesssim 1,$
since $k>1$. Using the above estimate, the substitution $\sigma\to-\sigma$ for the second summand and recalling the Maxwell operator notation (3.1), we obtain
${\langle}v{\rangle}^{k+1-\gamma}|Q_{sm}^{+}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(-\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v_{1}^{*})g_{k}(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\approx Q_{M,1}^{+}(|f_{l}|,|g_{k}|)+Q_{M,1}^{+}(|g_{k}|,|f_{l}|).$
Then, since $p<2$, Theorem 3.1 implies
(4.5) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{sm}^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L^{q}_{v}}\|g_{k}\|_{L^{p}_{v}}.$
Estimate for $Q_{r}^{+}$
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k+1-\gamma}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v|>4}\mathds{1}_{|v^{*}|\geq|v|/2}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim\frac{E^{\gamma/2}{\langle}v{\rangle}^{k+1-\gamma-l}}{(1+E)^{k/2}}\leq(1+E)^{\frac{1-l}{2}}\lesssim 1$
since $l>1$. Arguing as in the previous case, we obtain
(4.6) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{r}^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
It remains to estimate the singular operators $Q^{+}_{s,1}$ and $Q^{+}_{s,2}$.
We start with a preliminary observation: for $|v^{*}|<|v|/2$, we use (2.11), to write
$|v|\leq|v-v^{*}|+|v^{*}|\leq|R_{\sigma}^{+}(u)|+\frac{|v|}{2}\Rightarrow|v|\leq 2|R_{\sigma}^{+}(u)|=2|u||\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|.$
Now, since $0<\gamma\leq 1$ and $|u|\lesssim E^{1/2}$, we obtain
@@
(4.7) $\frac{|u|^{\gamma}{\langle}v{\rangle}^{k+1-\gamma}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{|v^{*}|<|v|/2}\lesssim\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{|v^{*}|<|v|/2}.$
Estimate for $Q_{s,1}^{+}$
Using (4.7), the exponents’ compatibility condition, and Hölder’s inequality, we obtain
@@
${\langle}v{\rangle}^{k+1-\gamma}|Q_{s,1}^{+}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}^{\frac{1}{q}-\frac{1}{p}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p}}\right)\left(\frac{{\langle}v{\rangle}^{2/p}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}$
(4.8) $\leq I_{1}^{1/q}(v)I_{2}^{1/p}(v),$